# Learning General-Purpose Biomedical Volume Representations using Randomized Synthesis

### v1

@@ 4.2 Few-shot 3D multi-label semantic segmentation
Data and setup. We evaluate few-shot segmentation performance on a diverse collection of datasets: cardiac MRI from MSD-Heart (Antonelli et al., 2022), abdominal CT from AMOS (Ji et al., 2022), prostate MRI from PROMISE12 (Litjens et al., 2014), abdominal MRI (Akin et al., 2016; Clark et al., 2013; Kavur et al., 2019; Linehan et al., 2016) from Learn2Reg-Abdomen (Hering et al., 2021), fetal brain MRI from FeTA (Payette et al., 2024), and an in-house dataset of whole uterus fetal BOLD MRI (WUFetal). WUFetal provides labels for the placenta, amniotic fluid, and the fetal brain and body. It is included as an out-of-distribution test dataset for the baselines below that are trained on multi-dataset collections of commonly imaged regions. Lastly, we operate in the few-shot regime, where only 1–3 annotated volumes per dataset are used for finetuning. The dataset splits are in App. B.5.
@@
Results. Table 2 and Fig. 6 present few-shot segmentation results. Our learning framework produces pretrained weights that consistently improve upon random initialization. Crucially, this improvement is achieved using only our dataset-agnostic pretrained weights and without any pretraining on unlabeled real volumes or data from similar domains, as used in previous work. Compared to larger foundation models specifically trained for segmentation, our pretrained general-purpose network achieves the first or second rank consistently and has the best average rank. Importantly, the second-best model changes dataset-to-dataset, indicating that the baseline methods do not generalize consistently to new biomedical contexts. We lastly emphasize that we achieve these gains with fewer parameters, without access to any real images, and while being applicable to other tasks as well, as described below.
@@ 4.3 Ablations and other analyses
Multitask capabilities. Current 3D biomedical vision foundation models are evaluated primarily using segmentation, raising the question of how well their features generalize to other tasks. To test this, we employ the few-shot segmentation baselines (that provide both pretrained encoders and decoders) as general-purpose feature extractors and use their features to drive a generic registration solver (ANTs), with implementation details provided in App. B.4.3. As reported in Table 2, all current 3D biomedical foundation models are unable to extract features that improve upon the baseline setting used by the solver (MutualInfo). In contrast, our network (Ours) outperforms it by wide margins and is the only method to yield multitask general-purpose features for the highly disparate tasks of multi-modality registration (Fig. 5) and few-shot segmentation (Table 2).
Label generation. To evaluate our label ensemble model, we replace it with other training labels while keeping the appearance model fixed. We test pretraining using: (a) synthetically generated labels that have no biomedical priors (Butoi et al., 2023; Hoffmann et al., 2021) (smshapes), (b) 1,573 label maps of real brain MRI using the FreeSurfer protocol (Fischl, 2012) (Brains) that represent real anatomical structures with dense per-voxel annotations, and (c) a combination of Brains and our model to mix real and synthetic sources of label maps. Details regarding these models are provided in App. B.6. Table 3 rows 1–4 show that both registration and segmentation performance decline with other choices of pretraining labels. Further, combining our labels with Brains does not affect segmentation but worsens abdominal registration. Lastly, pretraining on these alternative label models leads to unstable network features on real data (Fig. 7), likely explaining the performance degradation.
Temperature ($\tau$). The temperature hyperparameter $\tau$ in the contrastive loss (eq. 1) controls the penalty weight on negative pairs (Wang & Liu, 2021). In natural vision, smaller $\tau$ values (such as $\tau=0.07$ (Chen et al., 2020b)) are used to upweigh difficult negative pairs. However, as we train our network to learn similar representations within each label despite highly disparate appearances, the negative pairs are all difficult and require a relaxed $\tau$ of $0.33$ for stable training. We find that using lower $\tau$ leads to degenerate aliased features (Fig. 7) and worse registration results (Table 3). Interestingly, segmentation benefits slightly from an intermediate setting of $\tau=0.20$, indicating a tradeoff between optimal representations for downstream registration versus segmentation.
@@ 5 Discussion
Limitations and future work. Our approach does have limitations. We pretrained our network to be stable against intensity variations (among other variables) and demonstrated its utility for registration and segmentation. However, a small set of biomedical tasks rely on relative intensity values (Nakamura et al., 2017; Thomalla et al., 2011), making intensity invariance a suboptimal inductive bias for them. Further, while we operate on general volumetric tasks, certain biomedical inverse problems like MRI reconstruction take sensor-domain non-Cartesian (Schlemper et al., 2019) measurements as multichannel (Singh et al., 2022) inputs, requiring domain-specific architectural changes. Lastly, our segmentation experiments finetune our pretrained network on specific datasets, potentially introducing complexity for some clinical users. However, future extensions could directly use our proposed data engine to train promptable 3D segmentation models that require no such finetuning.
Conclusions. When combined with the right inductive biases, synthetic data models informed by biomedical templates enable the training of powerful general-purpose networks. This is important for 3D radiology, where existing annotated datasets are limited in sample size, often acquiring only dozens to at most a few thousand volumes. This leads to inflexible models that deteriorate under domain shifts. Trained only on synthetic volumes with our proposed framework, the resulting network provides substantial benefits on a variety of tasks. For example, its representations yield substantial improvements over the state-of-the-art in training-free multimodality deformable registration, a key area in biomedical vision. Further, the network can also serve as a downstream dataset-agnostic initialization for few-shot segmentation tasks and lead to improvements across multiple datasets.
@@ Acknowledgements
We gratefully acknowledge support from NIH NICHD R01HD100009, NIH NIBIB 1R01EB032708, NIH NIBIB R01EB033773, MIT-IBM Watson AI Lab, MIT Jameel Clinic, and the MIT CSAIL-Wistron Program.
@@ A.1 Additional synthetic data visualizations
Figure 9: Randomly selected synthetic volume (center slice) visualizations sampled from our proposed data engine (top). In the Brains and smshapes rows, we present samples corresponding to our ablations where we replace our proposed 3D label ensemble generator with real 3D brain labels (middle) or synthetically generated labels with no biomedical priors (bottom), respectively, but keep the appearance model unchanged.
@@ B.1.2 Volume synthesis
Figure 12: Post-Gaussian mixture model volume augmentation pipeline used to generate synthetic data for pretraining. Blue and red boxes refer to offline and online augmentations, respectively. All augmentations are applied to $128\times 128\times 128$ volumes. $p$ refers to the pre-defined probability of applying an individual augmentation and all other hyperparameters mentioned are consistent with MONAI (Cardoso et al., 2022) conventions.
@@ B.4.1 Baseline implementation
SynthMorph-shapes. SynthMorph-shapes is a domain-randomized diffeomorphic registration UNet trained on synthetic volume pairs generated from label maps. It is optimized using a Dice loss subject to diffusion regularization. We download its pretrained weights [^2] https://surfer.nmr.mgh.harvard.edu/ftp/data/voxelmorph/synthmorph/shapes-dice-vel-3-res-8-16-32-256f.h5 from the VoxelMorph library and use their Tensorflow-based registration framework. Lastly, as their network is fully convolutional, we use the input volumes at their native resolution without resizing.
@@
ConvexAdam. ConvexAdam is a high-performance multimodality registration solver and we use the b2671f8 commit of the repository [^4] https://github.com/multimodallearning/convexAdam/tree/b2671f86902390dec8dde702d0b583b451d84e98. Here we use the masked variant with default number of instance optimization iterations (80) and default hyperparameters of the MIND-SSC loss so as to use the same implementation of MIND-SSC across experiments. All the remaining parameters are fine-tuned on the validation data.
@@ B.5 Segmentation experiments implementation
Our UNet baselines use the crop sizes listed in Table 8 and our SwinUNETR-based pretrained baselines use the crop sizes used for pretraining and finetuning in their respective original papers for both consistency and due to their pretrained transformer backbones. All augmentations are performed online and the MRI and CT dataset experiments in Table 8 use the augmentations listed in Table 9 top and bottom, respectively. When finetuning, we train for 37,500 iterations with a batch size of four 3D crops using Adam with a starting step size of $2\times 10^{-4}$ cosine decayed to 0. Finally, we exclude the background label when reporting Dice statistics.
@@
Modality MRI MRI MRI MRI CT MRI

### v2

@@ 4.2 Few-shot 3D multi-label semantic segmentation
Data and setup. We evaluate few-shot segmentation performance on a diverse collection of datasets: cardiac bSSFP MRI from MSD-Heart (Antonelli et al., 2022), abdominal CT from AMOS (Ji et al., 2022), prostate T2w MRI from PROMISE12 (Litjens et al., 2014), abdominal SPIR MRI (Akin et al., 2016; Clark et al., 2013; Kavur et al., 2019; Linehan et al., 2016) from Learn2Reg-Abdomen (Hering et al., 2021), fetal brain HASTE MRI from FeTA (Payette et al., 2024), and an in-house dataset of whole uterus fetal BOLD MRI (WUFetal). WUFetal provides labels for the placenta, amniotic fluid, and the fetal brain and body. It is included as an out-of-distribution test dataset for our baselines below that have not been pretrained on fetal images. Lastly, we operate in the few-shot regime, where only 1–3 annotated volumes per dataset are used for finetuning. The dataset splits are in App. B.5.
@@
Results. Table 2 and Fig. 6 present few-shot segmentation results. Our framework produces pretrained weights that consistently improve upon random initialization. Crucially, this improvement is achieved using only our dataset-agnostic pretrained weights and without any pretraining on unlabeled real volumes or data from similar domains, as commonly done in previous work. Compared to larger foundation models specifically trained for segmentation, our pretrained general-purpose network achieves the first or second rank consistently and has the best average rank. Importantly, the second-best model changes dataset-to-dataset, indicating that the baseline methods do not generalize consistently to new biomedical contexts. We lastly emphasize that we achieve these gains with fewer parameters, without access to any real images, and while being applicable to other tasks as well, as described below.
@@ 4.3 Ablations and other analyses
Multitask capabilities. Current 3D biomedical vision foundation models are evaluated primarily for semantic segmentation, raising the question of how well their features generalize to other tasks. To test this, we utilize the few-shot segmentation baselines that provide both pretrained encoders and decoders as general-purpose feature extractors and use their features to drive a generic registration solver (ANTs), with implementation details provided in App. B.4.3. As reported in Table 2, all current 3D biomedical foundation models are unable to extract features that improve upon the baseline setting used by the default ANTs solver (MutualInfo). In contrast, our approach (Ours) outperforms it by wide margins and is thus the only method to yield multitask general-purpose features for the highly disparate tasks of multi-modality registration (Fig. 5) and few-shot segmentation (Table 2).
Label generation. To evaluate our label ensemble model, we replace it with other training labels while keeping the appearance model fixed. We test pretraining using: (a) synthetically generated labels that have no biomedical priors (Hoffmann et al., 2021) (smshapes), (b) 1,573 label maps of real brain MRI annotated using the FreeSurfer protocol (Fischl, 2012) (Brains) that represent real anatomical structures with dense per-voxel annotations, and (c) a combination of Brains and our model to mix real and synthetic sources of label maps. Details regarding these models are provided in App. B.6. Table 3 rows 1–4 show that both registration and segmentation performance decline with other choices of pretraining labels. Further, combining our labels with Brains does not affect segmentation but worsens abdominal registration. Lastly, pretraining on these alternative label models leads to unstable network features on real data (Fig. 7, right), likely explaining the performance degradation.
Temperature ($\tau$). The temperature hyperparameter $\tau$ in the contrastive loss (eq. 1) controls the penalty weight on negative pairs (Wang & Liu, 2021). In natural vision, smaller $\tau$ values (such as $\tau=0.07$ (Chen et al., 2020b)) are used to upweigh difficult negative pairs in the loss. However, as we train our network to learn similar representations within each label across highly disparate appearances, the negative pairs are all of higher difficulty and require a relaxed $\tau$ of $0.33$ for stable training. We find that using lower $\tau$ leads to degenerate aliased features (Fig. 7) and worse registration results (Table 3). Interestingly, segmentation benefits slightly from an intermediate setting of $\tau=0.20$, indicating a tradeoff between optimal representations for downstream registration versus segmentation.
@@ 5 Discussion
Limitations and future work. Our approach does have limitations. We pretrained our network to be stable against intensity variations (among other variables) and demonstrated its utility for registration and segmentation. However, a small set of biomedical tasks rely on relative intensities (Nakamura et al., 2017; Thomalla et al., 2011) and texture (Yu et al., 2017), making intensity invariance a suboptimal inductive bias for them. Further, while we operate on general volumetric tasks, certain inverse problems like MRI reconstruction take sensor-domain non-Cartesian (Schlemper et al., 2019) measurements as multichannel inputs, requiring domain-specific architectural changes (Singh et al., 2022). Lastly, our segmentation experiments finetune our pretrained network on specific datasets, potentially introducing complexity for some clinical users. However, future extensions could directly use our proposed data engine to train promptable 3D segmentation models that require no such finetuning.
Conclusions. When combined with the right inductive biases, synthetic data models informed by biomedical templates enable the training of powerful general-purpose networks. This is important for 3D radiology, where existing annotated datasets are limited in sample size, often acquiring only dozens to at most a few thousand volumes. This leads to inflexible models that deteriorate under domain shifts. Trained only on synthetic volumes with our proposed framework, the resulting network provides substantial benefits on a variety of radiological tasks. For example, its representations yield substantial improvements over the state-of-the-art in training-free multimodality deformable registration, a key area in biomedical vision. Further, the network can also serve as a downstream dataset-agnostic initialization for few-shot segmentation tasks and lead to improvements across multiple datasets.
@@ Acknowledgements
We gratefully acknowledge support from NIH NICHD R01HD100009, NIH NICHD 1R01HD114338, NIH NIBIB 1R01EB032708, NIH NIBIB R01EB033773, MIT-IBM Watson AI Lab, MIT Jameel Clinic, and the MIT CSAIL-Wistron Program.
@@ A.1 Additional synthetic data visualizations
Figure 9: Randomly selected synthetic volume (center slice) visualizations sampled from our proposed data engine (top). For clarity, these samples have not had the online augmentations in Figure 12 applied to them. In the Brains and smshapes rows, we present samples corresponding to our ablations where we replace our proposed 3D label ensemble generator with real 3D brain labels (middle) or synthetically generated labels with no biomedical priors (bottom), respectively, but keep the appearance model unchanged.
@@ A.5 Negative results
A.6 Additional results on ultrasound and microscopy datasets
Table 7: Few-shot segmentation Dice results on 3D ultrasound (SegThy) and 3D microscopy (NucMM-M) datasets, reported as mean Dice score with bootstrapped standard deviations. The best-performing method is in bold and the runner-up is underlined.
Method SegThy NucMM-M
Finetuning amount 2 subjects 1 volume
Number of classes 3 1
Random Init UNet .82(.03) .87(.03)
Transfer Learning .81(.02) .88(.02)
Models Genesis .78(.02) .83(.02)
MedicalNet .77(.03) .88(.03)
PrimGeoSeg .78(.03) .88(.03)
SMIT .78(.03) .91(.01)
Disruptive AE .74(.02) .88(.01)
Ours .84(.02) .89(.02)
The imaging modalities used in the datasets utilized for few-shot segmentation in Table 2 consist of CT and various MRI sequences. We extend our evaluation here to include two additional datasets representing distinct volumetric biomedical imaging modalities: 3D ultrasound and 3D microscopy.
[…]
In Table 7, for SegThy, the proposed model improves upon all baseline pretrained models by a clear margin, demonstrating our model’s usefulness to the unseen domain of 3D ultrasound. For NucMM-M, we achieve 2nd place performance to SMIT, albeit with low inter-method variability. However, as in Tables 22 and 7, our method consistently outperforms SMIT across all other seven segmentation tasks and both registration tasks, highlighting its generalizability.
A.7 Additional comparisons with 2D DINOv2
Table 8: Few-shot segmentation results (mean Dice score with bootstrapped standard deviations) comparing our method to slicewise 2D finetuning of DINOv2 (Oquab et al., 2023) finetuned with LoRA (Hu et al., 2021) and a trainable FPN decoder head and a randomly initialized 3D UNet.
MSD-Heart PROMISE12 L2RAb-MRI FeTA AMOS-CT WUFetal
2D DINOv2 w/ LoRA, FPN .78 (.01) .72 (.02) .72 (.07) .61 (.04) .25 (.01) .43 (.02)
3D UNet (Randomly Init.) .85 (.01) .80 (.02) .85 (.06) .78 (.03) .56 (.01) .73 (.02)
3D Ours .89(.01) .85(.01) .86(.06) .80(.03) .61(.01) .76(.02)
While our primary segmentation baselines are 3D foundation models specifically trained for radiology, recent work (Cekmeceli et al., 2024; Song et al., 2024) demonstrates that 2D natural vision foundation models, such as DINOv2 (Oquab et al., 2023), can be used for medical image analysis as well. In particular, concurrent work by Cekmeceli et al. (2024) shows that DINOv2 can be adapted for medical image segmentation in a fully supervised, slice-by-slice setting. To this end, we additionally benchmark against adapting DINOv2 for few-shot segmentation across all datasets in Table 2.
For DINOv2, we use the pretrained ViT-L/14 distilled (with registers) encoder. The encoder outputs an image embedding and we adapt it for 2D multi-label segmentation by attaching a trainable feature pyramid network (FPN) decoder head. We also train the encoder DINOv2 weights using low-rank adaptation (Hu et al., 2021) for efficiency. Since the encoder expects 3-channel inputs, single-channel medical images are repeated across channels. All other training details (e.g., training losses, iterations, learning rates) are matched to our 3D segmentation setup described in Appendix B.5 and we pick the weights with the best validation set performance to compute the test set scores below.
Table 8 shows that while the 2D foundation model can be finetuned slice-by-slice for medical image segmentation, our 3D biomedically-focused approach consistently outperforms it for few-shot segmentation across all six datasets and distinct application domains considered. We also find that training a well-tuned 3D UNet from random initialization generally outperforms slicewise application of a 2D natural image foundation model finetuned on that medical dataset, highlighting the need for natively 3D generalist approaches such as ours. This finding aligns with other works benchmarking 2D vs. 3D self-supervised segmentation approaches as well, as in Ren et al. (2022).
A.8 Additional comparisons with a abdominal CT foundation model trained on large-scale data
In parallel to our work, a large-scale annotated collection of 9,262 abdominal CT volumes was released (Li et al., 2024a) and is the basis for high-performance segmentation foundation models for abdominal CT (Li et al., 2024b). While these models perform well within their domain, it is unclear whether their pretrained weights transfer to new and unrelated biomedical domains and imaging modalities, such as fetal MRI. In contrast, our method, trained on highly variable synthetic data, does generalize consistently across diverse, unseen radiological domains and tasks.
To directly evaluate the transferability of abdominal CT foundation models to broader biomedical applications, we finetuned the SuPreM UNet (Li et al., 2024b), pretrained on AbdomenAtlas, on all non-abdominal CT datasets considered in the main paper. All splits, augmentations, and other training details are consistent with the main body of the paper.
Table 9: Few-shot segmentation results (mean Dice score with bootstrapped standard deviations) comparing our method with finetuning an abdominal CT pretrained segmentation model (SuPreM).
Methods MSD-Heart PROMISE12 L2RAb-MRI FeTA WUFetal
SuPreM .85 (.02) .75 (.02) .84 (.05) .79 (.03) .71 (.02)
Randomly Initialized UNet .85 (.01) .80 (.02) .85 (.06) .78 (.03) .73 (.02)
Ours .89 (.01) .85 (.01) .86 (.06) .80 (.03) .76 (.02)
Table 9 presents few-shot segmentation results for SuPreM, our model, and a randomly initialized UNet (with a matched architecture to ours), all finetuned on out-of-distribution datasets. SuPreM does not transfer well to new imaging modalities and biomedical applications. In contrast, our representations, trained only on highly variable synthetic data, generalize consistently.
We note that in future work, our framework can also be trained with supplemental real volumes and speculate that training on a mixture of synthetic and large-scale real datasets will have synergistic benefits, potentially yielding both high in-domain and out-of-domain performance.
@@ B.1.2 Volume synthesis
Figure 12: Post-Gaussian mixture model volume augmentation pipeline used to generate synthetic data for pretraining. Blue and red boxes refer to offline and online augmentations, respectively. All augmentations are applied to $128\times 128\times 128$ volumes. $p$ refers to the pre-defined probability of applying an individual augmentation and all other hyperparameters mentioned are consistent with MONAI (Cardoso et al., 2022) conventions. All augmentations use the default library hyperparameters besides Gamma Adjustments, Gaussian Blurs, and Gibbs Ringing, where we reduce the range of corruption to avoid degenerate samples.
@@ B.4.1 Baseline implementation
SynthMorph-shapes. SynthMorph-shapes is a domain-randomized diffeomorphic registration UNet trained on synthetic volume pairs generated from label maps. It is optimized using a Dice loss subject to diffusion regularization and has no tunable hyperparameters at test time. We download its pretrained weights [^2] https://surfer.nmr.mgh.harvard.edu/ftp/data/voxelmorph/synthmorph/shapes-dice-vel-3-res-8-16-32-256f.h5 from the VoxelMorph library and use their Tensorflow-based registration framework. Lastly, as their network is fully convolutional, we use the input volumes at their native resolution without resizing.
@@
ConvexAdam. ConvexAdam is a high-performance multimodality registration solver and we use the b2671f8 commit of the repository [^4] https://github.com/multimodallearning/convexAdam/tree/b2671f86902390dec8dde702d0b583b451d84e98. Here we use the masked variant with default number of instance optimization iterations (80) and default hyperparameters of the MIND-SSC loss so as to use the same implementation of MIND-SSC across experiments. All the remaining parameters are fine-tuned on the validation data, see Appendix A.2 for more details.
@@ B.5 Segmentation experiments implementation
Our UNet baselines use the crop sizes listed in Table 11 and our SwinUNETR-based pretrained baselines use the crop sizes used for pretraining and finetuning in their respective original papers for both consistency and due to their pretrained transformer backbones. All augmentations are performed online and the MRI and CT dataset experiments in Table 11 use the augmentations listed in Table 12 top and bottom, respectively. We found all baseline models and our model yielded better segmentation results when fully finetuned instead of other popular evaluation strategies for pretrained models such as linear probing. Additionally, we found that finetuning for a higher number of iterations with gradual learning rate decay consistently improved results on validation splits. Therefore, when finetuning, we train for 37,500 iterations with a batch size of four 3D crops using Adam with a starting step size of $2\times 10^{-4}$ cosine decayed to 0. Finally, we exclude the background label when reporting Dice statistics.
@@
Modality BOLD MRI T2w MRI bSSFP MRI SPIR MRI CT HASTE MRI